When light passes through a diamond, the speed of light is 1/2.42 times the speed of light in vacuum. The frequency of yellow light is 5.3 x 1014 Hz. The frequency doesn’t change as the yellow light passes through a diamond. What is the wavelength of yellow light when it is passing through diamond?

The speed of light in a vacuum is 3x108 m/s.
1
Expert's answer
2020-01-24T05:17:47-0500
"\lambda=\frac{c}{nf}"

"\lambda=\frac{3\cdot 10^8}{(2.42)5.3\cdot 10^{14}}=2.34\cdot 10^{-7}m=234\ nm"